OpenStudy (anonymous):

what are the next two numbers in the pattern bellow 7,21,63,189..... Medals Will Be Given

OpenStudy (anonymous):

pattern is multiply by 3 189 * 3 =567 567*3 = 1701

OpenStudy (anonymous):

it's a GP with r= 3

OpenStudy (anonymous):

an=a1*(r^(n-1))